OpenStudy (anonymous):

pls help... log6(x+9)-log36x=1

OpenStudy (zepp):

Use the property of \[\large \log(a)-\log(b)=\log\frac{a}{b}\]

OpenStudy (zepp):

\[\large \log(6(x+9))-\log(36x)=\log\frac{6(x+9)}{(36x)}\] I'm pretty sure you can continue all by yourself now! :D

OpenStudy (anonymous):

is it x+9/6x